What is the sum of the roots of 6x^2 - x - 1 = 0
Answer by TimothyLamb(4379)

6x^2 - x - 1 = 0
---
the above quadratic equation is in standard form, with a=6, b=-1, and c=-1
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
6 -1 -1
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
x = 0.5
x = -1/3
---
answer:
sum of the root = 1/2 - 1/3
3/6 - 2/6
1/6
0.166666
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php